OpenStudy (harrycraig):

IF the roots of the equation x^2-4x-log3a=0 (log 3base a ) are real , then the least value of a

OpenStudy (harrycraig):

\[x^2-4x-\log3^a=0\]

OpenStudy (harrycraig):

@pooja195

jimthompson5910 (jim_thompson5910):

I'm going to replace the 'a' with y hint: if the two roots are real, then D = b^2 - 4ac is going to be greater than or equal to zero. Ie, b^2 - 4ac >= 0 in this case, a = 1, b = -4, c = log(3,y)

OpenStudy (harrycraig):

i am getting problem with log case . the way of solving i know .@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large {\color{red}{1}}x^2+({\color{green}{-4}}x)+{\color{blue}{\log_{3}(y)}}\] is in the form \[\Large {\color{red}{a}}x^2+{\color{green}{b}}x+{\color{blue}{c}}\] where \[\Large \color{red}{a = 1}\] \[\Large \color{green}{b = -4}\] \[\Large \color{blue}{c = \log_{3}(y)}\]

jimthompson5910 (jim_thompson5910):

\[\Large D \ge 0\] \[\Large b^2 - 4ac \ge 0\] \[\Large (-4)^2 - 4(1)(\log_3(y)) \ge 0\] \[\Large 16 - 4*\log_3(y) \ge 0\] solve for y

OpenStudy (harrycraig):

how to exclude log that's my problem

jimthompson5910 (jim_thompson5910):

\[\Large 16 - 4*\log_3(y) \ge 0\] \[\Large 16 \ge 4*\log_3(y)\] \[\Large 4*\log_3(y) \le 16\] \[\Large \log_3(y) \le 4\] \[\Large y \le 3^4\] \[\Large y \le 81\] So as long as y is less than or equal to 81, then \(\Large x^2 - 4x + \log_{3}(y) = 0\) will have two real solutions for x.

jimthompson5910 (jim_thompson5910):

oh wait, it's not +log(3,y) it's -log(3,y) so that will change things a bit

jimthompson5910 (jim_thompson5910):

a = 1 b = -4 c = -log(3,y) \[\Large D \ge 0\] \[\Large b^2 - 4ac \ge 0\] \[\Large (-4)^2 - 4(1)(-\log_3(y)) \ge 0\] \[\Large 16 + 4*\log_3(y) \ge 0\] solve for y

OpenStudy (harrycraig):

yeah it changed the whole answer

jimthompson5910 (jim_thompson5910):

The steps will be very similar

jimthompson5910 (jim_thompson5910):

first try to isolate the log itself, then isolate y after that

OpenStudy (harrycraig):

it asked the least value so will the answer be 81 or 1/81

jimthompson5910 (jim_thompson5910):

were you able to solve \[\Large 16 + 4*\log_3(y) \ge 0\] for y?

OpenStudy (harrycraig):

y<=81 thanks

jimthompson5910 (jim_thompson5910):

no

OpenStudy (harrycraig):

why

jimthompson5910 (jim_thompson5910):

\[\Large 16 + 4*\log_3(y) \ge 0\] \[\Large 4*\log_3(y) \ge -16\] \[\Large \log_3(y) \ge -4\] \[\Large y \ge 3^{-4}\] \[\Large y \ge \frac{1}{3^4}\] \[\Large y \ge \frac{1}{81}\]

jimthompson5910 (jim_thompson5910):

So as long as y is greater than or equal to \(\Large \frac{1}{81}\), then \(\Large x^2 - 4x - \log_{3}(y) = 0\) will have two real solutions for x.

OpenStudy (harrycraig):

thanks

jimthompson5910 (jim_thompson5910):

np